How much would $120 invested at 6% interest compounded annually be worth after 21 years? Round your answer to the nearest cent. Do not include units in your answer.

Answer by rfer(10432)

FV=P(1+r)^n
FV=120(1.06)^21
FV=120(3.3995636)
FV=407.95